### v2

# Text and Vision [Version 0.2: For Nano-Banana Like Generation]

Abstract
Chain-of-Thought (CoT) reasoning has been widely adopted to enhance the performance of Large Language Models (LLMs) on complex tasks by step-by-step problem solving. However, extending CoT to complex vision-language tasks remains challenging, as it often requires interpreting visual change to assist reasoning. Existing methods often struggle with this due to limited capacity of modeling visual state transitions or incoherent visual trajectories caused by fragmented architectures.
To overcome these limitations, we propose Uni-CoT, a Unified Chain-of-Thought framework that enables coherent visual state transition and grounded text reasoning for multi-modal tasks. Our key idea is to leverage a unified model, capable of both image understanding and generation, to reason over visual content and model visual transitions. However, empowering a unified model to achieve that is non-trivial due to the high computational cost and instable training. To address this, Uni-CoT introduces a novel two-level reasoning paradigm: A Macro-Level CoT for high-level task planning, and A Micro-Level CoT for subtask execution. This hierarchical design significantly reduces computational overhead by localizing reasoning dependencies. Furthermore, we introduce a structured training paradigm that combines multi-modal supervision for macro-level CoT with multi-task supervision for micro-level CoT. Together, these innovations allow Uni-CoT to perform scalable and stable multi-modal reasoning with coherent visual transitions. Furthermore, with our designs, our model can be efficiently trained only by 8 A100 GPUs. Experimental results on reasoning-driven image generation benchmark (WISE) and editing benchmarks (RISE and KRIS) indicates that Uni-CoT demonstrates state-of-the-art performance and strong generalization, establishing Uni-CoT as a promising solution for multi-modal reasoning. Project Page and Code: https://sais-fuxi.github.io/projects/uni-cot/
Figure 1: The multi-modal reasoning trajectory of Uni-CoT. Uni-CoT extends Chain-of-Thought to the multi-modal domain, enabling a unified model to perform coherent, grounded, and step-by-step reasoning across text and images. More results refer to Figure S3.
* Equal contribution, # project leader, $\dagger$ Corresponding authors.Work in progress.
1 Introduction
Chain-of-Thought (CoT) [68] is an approach that enhances the performance of Large Language Models (LLMs) on complex tasks by imitating the way humans solve problems. It works by encouraging LLMs to produce explicit intermediate reasoning steps, allowing them to tackle complex tasks in a step-by-step manner. Motivated by its success [20, 14, 41], recent works [84, 42, 87] have paid efforts to extend CoT to multi-modal domain, aiming to equip Multi-modal Large Language Models (MLLMs) with the reasoning capacity to handle complex vision-language tasks, such as complex visual question answering [75], image editing [21], and embodied planning [43].
Prior efforts [82, 78, 19, 30, 72, 65, 57] have employed reinforcement learning (RL) to enhance the text-based reasoning capabilities of multi-modal large language models (MLLMs). While effective in text-centric applications, these approaches reveal notable limitations in domains that require strong visual reasoning, such as geometric analysis or visual navigation [70, 12]. In fact, many of these visual reasoning tasks can be easily solved by middle-school students, underscoring a fundamental gap between MLLMs and human cognitive abilities. Existing studies [17, 60] suggest that this human advantage stems from the natural integration of visual state transitions into reasoning. Such skills are cultivated early in multi-modal environments, where even simple activities like cooking require interpreting textual instructions while monitoring visual cues. However, this capacity remains challenging for current MLLMs to replicate, highlighting the need for future models to explicitly embed visual state transitions into their reasoning processes to unlock more general multi-modal reasoning abilities.
Motivated by this gap, recent studies [27, 58, 28, 29, 55, 22] have explored programmatic visual manipulations (e.g., cropping, line plotting) to approximate visual state transitions. While effective at capturing local image dynamics, these methods fall short in modeling the structural visual changes needed for complex tasks such as goal-directed navigation or object rearrangement. To overcome this, other approaches [76, 86, 25] couple MLLMs with image or video generators, thereby enabling larger-scale transitions. In parallel, recent works [26, 32] leverage external reward models to guide generators through iterative visual refinements. However, the loose integration between reasoning and generation in these frameworks often results in fragmented reasoning flows and inconsistent transitions, ultimately limiting coherent multi-modal reasoning.
To address these limitations, we propose Uni-CoT, a Unified Chain-of-Thought framework, designed to support structural visual transitions and coherent text understanding for multi-modal tasks. Our preliminary idea is to leverage a unified model [15, 59, 74, 11, 61], capable of both image understanding and generation, to support reasoning grounded in visual content and modeling dynamic visual state transitions. This design is motivated by the intuition that using a single model for both reasoning and generation naturally minimize the discrepancy between reasoning trajectories and visual transitions, thereby ensuring coherence in multi-modal reasoning. In addition, a unified architecture enables seamless implementation of end-to-end fine-tuning and reinforcement learning, facilitating higher coherence and performance ceilings for multi-modal tasks.
However, enabling a unified model to perform multi-modal reasoning is non-trivial due to two major challenges: (1) Computational Burden: Unlike text-only reasoning which consumes roughly= 300 tokens per step, multi-modal reasoning need to jointly produce textual and visual intermediates, requiring up to 10,000 tokens per step (see Sec. 2). This dramatic increase in sequence length greatly amplifies both computational and storage overhead. (2) Training Instability: Longer sequences demand the ability to model long-range dependencies, making optimization more difficult. In addition, interleaved image-text generation exacerbates instability, as the two modalities exhibit mismatched learning dynamics. Reinforcement learning further compounds these issues, requiring carefully designed modality-aware rewards and optimization strategies.
Fortunately, humans face similar difficulties when solving multi-modal complex tasks and offer insights into potential solutions. In real-world, complex multi-modal reasoning places heavy demands on working memory and often becomes challenging when integrating different types of information with multiple input channels [47]. For instance, solving a large jigsaw puzzle through brute-force trial quickly becomes intractable as the search space grows exponentially, overwhelming cognitive capacity. A well-established strategy for overcoming such challenges is the hierarchical organization of reasoning [5, 45]. At a higher level, abstract reasoning generates a global plan that decomposes the problem into manageable subtasks; at a lower level, situated reasoning focuses on executing these subtasks with minimal interference. Returning to the puzzle example, experienced solvers often segment the board into smaller regions based on color or object similarity as a high-level planning and then assemble small areas one by one as low-level inference. The successful assembly of local areas provides intermediate feedback, progressively guiding the completion of the entire puzzle. This dual-level structure shortens reasoning trajectories and reduces cognitive load, enabling humans to sustain coherent reasoning over long horizons without being overwhelmed by complexity.
Inspired by this human cognitive ability, we propose a macro–micro hierarchical reasoning framework to mitigate the complexity of multi-modal reasoning. At the macro-level CoT, our model performs global planning by decomposing a complex task into manageable subtasks and synthesize the final result only according to subtasks’ results. At the micro-level CoT, the model focuses on solving each subtask with minimal interference from irrelevant context. To minimize the computational burden, the micro-level CoT is further formulated as a Markov Decision Process (MDP). By transforming a lengthy and cognitively demanding reasoning trajectory into a series of structured reasoning blocks, our hierarchical design reduces redundant token interactions in CoT, thereby substantially lowering computational complexity and improving training stability.
Building on this hierarchical design, we further decompose multi-modal CoT learning into two complementary components: (1) Macro-Level CoT Modeling. The model is directly refined on interleaved text–image content to acquire the ability of global planning and final result synthesis. (2) Micro-Level CoT Modeling. The micro-level reasoning is further decomposed into an MDP-style process, where we design four auxiliary tasks, such as action generation and reward estimation, to facilitate effective and efficient learning. This decoupled paradigm provides supervision at both global and local levels, thereby enabling scalable and efficient training for complex multi-modal reasoning tasks. Moreover, to enhance optimization stability, we introduce a node-based reinforcement learning strategy. With these training designs, Uni-CoT can be trained efficiently on only 8 A100 GPUs. Extensive experiments on reasoning-driven image generation and editing benchmarks show that Uni-CoT delivers state-of-the-art performance, establishing it as a promising solution for multi-modal reasoning.
2 Preliminary: The Unifed Model - BAGEL
To enable unified chain-of-thought (CoT) reasoning across both image and text modalities, our framework builds upon BAGEL (Scalable Generative Cognitive Model), a recently released open-source unified model that supports joint vision-language understanding and generation. We briefly introduce the core elements of BAGEL in below.
Architecture.
BAGEL adopts a unified decoder-only Transformer architecture and incorporates a Mixture-of-Transformer-Experts [36] design. It is initialized from Qwen-2.5 [2]. Bagel consists of two experts: one dedicated to understanding and the other to generation. Both experts operate over shared multi-modal token sequences through a unified self-attention mechanism, allowing flexible and lossless fusion across modalities without introducing task-specific bottlenecks. Specifically, BAGEL integrates two modality-specific visual encoders:
Vision Transformer (ViT) encoder, initialized from SigLIP2 [62], for image understanding. It transforms an image into approximately 4,900 tokens. Variational Autoencoder (VAE), initialized from FLUX [35], for image generation. It encodes an image into a latent grid of $\times64$, resulting in 4,096 tokens.
With these two visual encoders, BAGEL employs an expert-routing mechanism with hard gating: the understanding expert activates the ViT encoder for image understanding and text generation, while the generation expert activates the VAE decoder for high-quality image synthesis conditioned on text and images. This dual-pathway design enables fine-grained visual generation and high-level semantic grounding within a unified autoregressive modeling framework.
Training Objectives.
BAGEL is jointly trained for multi-modal understanding and generation through two complementary loss functions:
  • Cross-Entropy Loss for text prediction:
    $\mathcal{L}_{\text{CE}}^{\text{text}}=\sum_{i=1}^{C}x_{i}\log(\hat{x}_{i}),$ (1)
    where $x_{i}$ denotes the target text token, $\hat{x}_{i}$ is the predicted token and $C$ is the number of token classes.
  • Mean Squared Error (MSE) Loss. For denoising-based generation under the Rectified Flow [40] paradigm, given a clean latent $\mathbf{x}_{0}$ and a Gaussian noise sample $\mathbf{x}_{1}$, we construct the noisy latent $\mathbf{x}_{t}$ via linear interpolation:
    $\mathbf{x}_{t}=(1-t)\cdot\mathbf{x}_{0}+t\cdot\mathbf{x}_{1},\quad t\in[0,1]$ (2)
    The model is then trained to predict the velocity field using the MSE loss:
    $\mathcal{L}_{\text{MSE}}^{\text{image}}=\mathbb{E}\left[\left\|\mathbf{g}_{\theta}(\mathbf{x}_{t}\mid\mathbf{c})-\left(\mathbf{x}_{0}-\mathbf{x}_{1}\right)\right\|^{2}\right]$ (3)
    Here, $\mathbf{g}_{\theta}(\mathbf{x}_{t}\mid\mathbf{c})$ denotes the velocity predicted by the model $\mathbf{g}$, conditioned on noisy latent $\mathbf{x}_{t}$ and context $\mathbf{c}$.
  • The total loss is a combination of both objectives:
    $\mathcal{L}_{\text{total}}=\lambda_{\text{CE}}\cdot\mathcal{L}_{\text{CE}}^{\text{text}}+\mathcal{L}_{\text{MSE}}^{\text{image}},$ (4)
    where ${\lambda}_{\text{CE}}$ is a coefficient to balance two losses.
Inference.
At inference time, BAGEL operates autoregressively over interleaved multi-modal token sequences: Understanding tasks (e.g., visual question answering, multi-modal reasoning): BAGEL consumes ViT-encoded image tokens and text tokens, and outputs the next text token via standard next-token prediction. Generation tasks (e.g., text-to-image synthesis, image editing): BAGEL predicts VAE tokens conditioned on text or image prompts via Rectified Flow process. This unified inference mechanism enables BAGEL to handle diverse multi-modal tasks within a single interface.
High Complexity for Multi-Modal Reasoning in BAGEL.
As discussed above, while BAGEL provides a unified modeling backbone, it faces significant computational bottlenecks when applied to step-wise multi-modal reasoning. Unlike text-only CoT, where each reasoning step typically involves 300 tokens, multi-modal CoT often requires both image understanding and image generation within each step, introducing substantial overhead: generating an image via VAE incurs approximately 4,096 tokens and encoding an image via ViT for understanding introduces an additional 4,900 tokens. This results in nearly 9,000 visual tokens, in addition to around 300 text tokens, per reasoning step, leading the overall training and inference costs to be prohibitively expensive. Moreover, this token-intensive formulation imposes significant challenges on model optimization, often hindering convergence and limiting generalization, particularly in complex tasks that require long and compositional reasoning chains.
3 Method
Figure 2: Overview of the Uni-CoT framework. Uni-CoT consists of two complementary reasoning branches: (1) Macro-Level CoT, which decomposes a complex task into simpler subtasks and aggregates their outcomes to synthesize the final result. To reduce learning and computational overhead, intra-subtask reasoning is kept implicit. This process is enforced through a macro attention mask that reveals only the system prompt, high-level plans, and subtask outputs. (2) Micro-Level CoT, which executes individual subtasks while filtering out irrelevant information. It is modeled as a Markov Decision Process (MDP), where each reasoning and self-reflection step depends solely on the previous state and the current prompt. This process is enforced through a micro attention mask that restricts visibility to the last state and current instruction. High-resolution depictions of the macro and micro attention masks are shown in Figure S1.
To tackle the inherent complexity of multi-modal CoT, Uni-CoT adopts a hierarchical framework to minimize the token interactions in reasoning. As illustrated in Figure 2, the reasoning process is organized into two levels: a Macro-Level CoT, responsible for high-level task planning and summarization (Sec.3.1); and a Micro-Level CoT, which generates stable and reliable results for each subtask (Sec.3.2). This two-tiered design enables more efficient reasoning, stronger generalization, and improved interpretability across diverse multi-modal tasks. The training procedures of this framework are further detailed in Sec. 3.3.
3.1 Macro-Level CoT: Planning Strategies
Given a complex task, humans typically begin by outlining one or more abstract pathways toward the goal, rather than considering low-level details from the outset [13, 17]. Inspired by this cognitive behavior, a core function of the Macro-Level CoT is to formulate a global plan that decomposes the task into a set of simple, tractable subtasks.
Specifically, drawing inspiration from [16, 77, 49, 79], we propose three global planning mechanism:
1. Sequential Decomposition: Humans often tackle complex tasks by addressing intermediate goals in a step-by-step manner. Analogously, we define a sequential decomposition strategy, wherein a task is split into a sequence of subtasks, to simplify the overall reasoning path and improve traceability. 2. Parallel Decomposition: Humans often decompose complex tasks into several independent components that can be solved concurrently, thereby accelerating the overall process. Inspired by this strategy, we propose a parallel decomposition approach that allows the model to reason over multiple subtasks simultaneously, improving both efficiency and scalability. 3. Progressive Refinement: In uncertain environments, such as navigating a maze, human often forego rigid plans and instead refine their decisions iteratively. To emulate this behavior, we propose a progressive refinement strategy, wherein the model incrementally refines its plan and revise earlier steps if inconsistencies arise, thereby supporting adaptive reasoning. These strategies empower the Macro-Level CoT to effectively address the “what-to-do” aspect of reasoning.
In addition to high-level planning, the Macro-Level CoT is responsible for analyzing and synthesizing the outputs of individual subtasks into a coherent final solution, as illustrated in Figure 2. Importantly, the internal reasoning processes within each subtask are abstracted away from the macro perspective, enabling the macro planner to concentrate solely on structural decomposition and the global reasoning trajectory. To enforce this separation of concerns during both training and inference, we introduce a macro attention mask (Figure 2 or Figure S1 (a)). This mechanism selectively exposes only the system prompt, macro-level planning outputs, the outcome of each subtask, and the final result. All intermediate reasoning traces, such as textual rationales or image edits generated during micro-level execution, are fully masked. By restricting visibility to high-level signals, the Macro-Level CoT preserves a top-down reasoning perspective, thereby enhancing modularity and scalability in multi-modal CoT.
Furthermore, with these designs, training the model to learn the Macro-Level CoT is simplified into two key tasks: 1) generating global plans based on user text and image prompts, and 2) synthesizing the final results based on the outputs of the subtasks. These focused training objectives significantly stabilize and simplify the training process.
3.2 Micro-Level CoT: Subtask Execution
Once a subtask is assigned by the Macro-Level planner, the Micro-Level CoT is responsible for its execution. Since the overall reliability of a system hinges on the consistency and coherence of outputs across subtasks, the core objective of the Micro-Level CoT is to generate high-reliable and high-quality results for the assigned subtask.
Figure 3: MDP-based reasoning architecture. (a) Overview of the sequential MDP process for multi-modal reasoning. (b) Architecture of a single MDP step $(s_{t},a_{t},s_{t+1},r_{t+1})$. The transition from one state to the next is guided by the subtask instruction, with the learnable content highlighted in pink.
To support this, we introduce a Self-Reflection mechanism to enhance both robustness and adaptability for subtask execution. Specifically, as illustrated in Figure 2, after attempting to complete a subtask, the model evaluates the quality of its output and determines whether revision is necessary. If logical inconsistencies or cross-modal mismatches are detected, the model revises its output and re-evaluates the result in a closed-loop feedback cycle. This process continues until the model deems the output satisfactory, thereby concluding the subtask execution. This mechanism ensures that generated outputs are aligned with the model’s internal reasoning and multi-modal understanding.
To further reduce the complexity of modeling the self-reflection mechanism, we treat this reasoning process as a Markov Decision Process (MDP) [51], where each neighboring reasoning and self-reflection sequence pair depends only on the outcome of the previous step and the given subtask instruction. As shown in Figure 3 (a), we define the current reasoning result (a multi-modal pair) as MDP node $t$, with the revised output after self-evaluation and editing as MDP node $t+1$. We assume that the transition from node $t$ to $t+1$ is solely governed by the node $t$ and the fixed subtask instruction, mirroring human behavior during self-reflection, where attention is focused on assessing the current state. To enforce this locality, we introduce a micro attention mask, which limits attention to short-range state transitions, as illustrated in Figure 2. This approach simplifies the learning objective while enhancing both training stability and computational efficiency.
Formally, as illustrated in Figure 3 (b), each element in our MDP step $(s_{t},a_{t},s_{t+1},r_{t+1})$ is defined as follows: State $s_{t}$: The current state that consists of the previous step’s results, including both textual and visual content; Action $a_{t}$: A hybrid operation that consists of textual editing prompt generation $a^{text}_{t}$ and corresponding image editing $a^{image}_{t}$; Next state $s_{t+1}$: The updated state, encompassing both the edited image and an aligned textual summary; Reward $r_{t+1}$: A textual judgment measuring alignment between the next state’s result and the subtask objective. With this formulation, the learning objectives of the Micro-Level CoT are formulated into two core competencies: (1) subtask completion, which involves both image editing and multi-modal understanding; and (2) MDP modeling, including hybrid action generation, next-state prediction, and reward estimation, highlighted in pink in Figure 3 (b).
3.3 Training Paradigm
The Uni-CoT training is conducted in two sequential stages: Supervised Fine-Tuning and Reinforcement Learning.
Supervised Fine-Tuning (SFT).
As detailed in above sections, SFT comprises two complementary components: Macro-Level CoT learning and Micro-Level CoT learning.
For the Macro-Level CoT, we adopt a joint loss that supervises both text and image generation for planning and final synthesis. Specifically, following Eq. 4, we apply cross-entropy (CE) loss for textual output and mean squared error (MSE) loss for image generation:
$\mathcal{L}_{\text{Macro}}={\lambda}_{\text{CE}}\cdot\mathcal{L}_{\text{CE}}^{\text{text}}+\mathcal{L}_{\text{MSE}}^{\text{image}},$ (5)
where ${\lambda}_{\text{CE}}$ is a balancing coefficient that controls the relative importance of textual and visual losses.
For the Micro-Level CoT, subtask completion is supervised using interleaved multi-modal data, similarly adopting CE and MSE losses for model learning. Additionally, to learn MDP-based self-reflective reasoning process, we decompose learning into four auxiliary objectives: text action $a^{text}_{t}$ generation, image action generation, next-state prediction and reward estimation. More details and data structures are are provided in Appendix Sec. B.
Reinforcement Learning (RL).
To further enhance reasoning robustness and adaptability, we adopt reinforcement learning (RL) to optimize both Macro-Level and Micro-Level reasoning behaviors. Rewards are designed to reflect subtask completion quality and overall task success.
In this work, we employ a simplified yet effective RL strategy, Direct Preference Optimization (DPO) [53, 63], to align model outputs with human-preferred reasoning trajectories. We decouple the DPO training into two stages:
  • Textual Preference Learning: Encourages the model to prefer coherent and correct reasoning paths over suboptimal or inconsistent alternatives via pairwise annotations.
  • Visual Preference Learning: Guides image editing behavior by optimizing toward preferred visual outputs, based on human or proxy feedback that captures semantic correctness and visual fidelity.
Further details on the reinforcement learning setup will be provided in a future version of this work.
4 Experiments
Figure 4: Qualitative Results for Reliable Image Generation. Uni-CoT demonstrates impressive image generation capabilities on complex, abstract, and reasoning-intensive prompts. Notably, these results are achieved through joint image-text reasoning, where Uni-CoT iteratively evaluates the current visual state, provides textual instructions for modification, and then executes those modifications.
4.1 Implementation Details
Dataset.
We construct a multi-modal reasoning dataset to support both macro- and micro-level CoT learning for our Uni-CoT model. At the macro level, seed prompts are augmented with explicit logical deductions and enriched visual details, then decomposed into 2–3 subtasks using large language models. Each subtask yields interleaved text–image traces comprising planning, evaluation, refinement instructions, and intermediate generations. At the micro level, preliminary images are generated and refined through iterative self-reflection loops, where models repeatedly evaluate, adjust, and edit outputs to capture fine-grained reasoning and editing trajectories.
Through this pipeline, our dataset comprises approximately 11K interleaved text–image pairs for macro-level planning traces and 20K pairs for micro-level refinement loops, spanning both text and image modalities. We also crop several data from [4, 67, 8, 80] to enhance the generalization of our model across various scenarios. Additional details on dataset construction and statistics are provided in Appendix Sec. C.
Training Details.
Owing to our decomposition of interleaved image-text CoT via MDP modeling and high-quality data construction, the training of Uni-CoT is quite efficient, as all experiments can be accomplished on 8 NVIDIA A100 GPUs. Following common practices, we utilize FlashAttention, Fully Sharded Data Parallel (FSDP), and mixed-precision training for better computation efficiency. Throughout the training process, all the parameters in the unified model are optimized using Adam optimizer with a constant learning rate of 2e-5. Additionally, we employ a linear warmup learning rate schedule, increasing the learning rate from zero over the first 200 steps. During each training step, we combine all data used for understanding and generation expert training, randomly sample from the combined dataset, and pack the samples into sequences with a target length of 32,768 tokens.
Table 1: Quantitative evaluation results on GenEval [24]. Notably, ‘Gen. Only’ stands for an image generation model, and ‘Unified’ denotes a model that has both understanding and generation capabilities. $\dagger$} refers to the best preformance of Bagel reproduced by ourselves. Uni-CoT Init is our initial results without multi-modal reasoning.
[…]
Gen. Only PixArt-$\alpha$ [9] 0.98 0.50 0.44 0.80 0.08 0.07 0.48
DALL-E $2$ [54] 0.94 0.66 0.49 0.77 0.10 0.19 0.52
Emu33-Gen  [66] 0.98 0.71 0.34 0.81 0.17 0.21 0.54
SDXL [50] 0.98 0.74 0.39 0.85 0.15 0.23 0.55
DALL-E $3$ [3] 0.96 0.87 0.47 0.83 0.43 0.45 0.67
SD3-Medium [18] 0.99 0.94 0.72 0.89 0.33 0.60 0.74
FLUX.1-dev [34] 0.98 0.93 0.75 0.93 0.68 0.65 0.82
Unified LWM [37] 0.93 0.41 0.46 0.79 0.09 0.15 0.47
SEED-X [23] 0.97 0.58 0.26 0.80 0.19 0.14 0.49
TokenFlow-XL [52] 0.95 0.60 0.41 0.81 0.16 0.24 0.55
ILLUME [64] 0.99 0.86 0.45 0.71 0.39 0.28 0.61
Emu33-Gen [66] 0.99 0.81 0.42 0.80 0.49 0.45 0.66
Show-o [74] 0.98 0.80 0.66 0.84 0.31 0.50 0.68
Janus-Pro-7B [10] 0.99 0.89 0.59 0.90 0.79 0.66 0.80
MetaQuery-XL [48] - - - - - - 0.80
BAGEL [15] 0.99 0.92 0.78 0.87 0.53 0.64 0.79
Uni-CoT 0.99 0.95 0.82 0.89 0.60 0.72 0.83
Uni-CoT Init 0.99 0.95 0.82 0.90 0.55 0.69 0.81
Uni-CoT 0.99 0.96 0.84 0.92 0.57 0.71 0.83
Table 2: Quantitative evaluation results on WISE [46]. Notably, the evaluation results are averaged over five independent runs to ensure statistical robustness. As can be drawn, Uni-CoT achieves the best performance among its open-source alternatives. Uni-CoT Init is our initial results without multi-modal reasoning.
Model Culture↑ Time↑ Space↑ Biology↑ Physics↑ Chemistry↑ Overall↑
Janus [69] 0.16 0.26 0.35 0.28 0.30 0.14 0.23
MetaQuery [48] 0.56 0.55 0.62 0.49 0.63 0.41 0.55
Bagel-Think [15] 0.76 0.69 0.75 0.65 0.75 0.58 0.70
Uni-CoT 0.76 0.70 0.76 0.73 0.81 0.73 0.75
GPT-4o [31] 0.81 0.71 0.89 0.83 0.79 0.74 0.80
Uni-CoT Init 0.75 0.67 0.79 0.64 0.79 0.65 0.72
Uni-CoT 0.76 0.70 0.76 0.73 0.81 0.73 0.75
4.2 Experimental Setup
In this work, we focus on the Uni-CoT on two generation tasks in main paper: image generation and image editing.
For the image generation task, following [15], we conduct experiments on two widely adopted benchmarks: GenEval [24] and WISE [46]. GenEval [24] serves as a general benchmark for assessing object-focused text-to-image alignment, while WISE [46] is a reasoning-driven benchmark designed to evaluate a model’s ability to generate faithful outputs from abstract, reasoning-intensive prompts.
For the image editing task, we benchmark Uni-CoT on GEdit-Bench [38], RISE [85], and KRIS [71]. GEdit-Bench [38] is a general benchmark that provides a diverse set of real-world editing tasks. In contrast, RISE [85] focuses on reasoning-informed editing across temporal, causal, spatial, and logical dimensions, while KRIS [71] serves as a diagnostic benchmark categorizing editing tasks into factual, conceptual, and procedural knowledge types.
4.3 Experimental Setup
Quantitative Results.
Table 1 presents the evaluation results of Uni-CoT and other baselines on GenEval [24], which benchmarks basic image generation capabilities. Among the compared models, Uni-CoT outperforms its base model Bagel. We attribute this improvement primarily to the self-reflection mechanism, which effectively corrects errors in the initial generations.
Table 2 reports the evaluation results of Uni-CoT and other baselines on WISE, benchmarking their reasoning-based image generation ability across multiple knowledge domains. Among the evaluated models, Uni-CoT consistently achieves state-of-the-art results across all the evaluated domains, demonstrating significantly better reasoning-based image generation capability compared to open-source baselines.
Qualitative Analysis.
As shown in Figure 4, Uni-CoT is capable of correcting semantically inaccurate generations via multi-round self-reflection. For prompts requires complex reasoning over commonsense and their visual appearance, the initial outputs of Uni-CoT may appear plausible but deviate from the intended semantics. However, by employing a multi-round self-reflection mechanism, Uni-CoT can iteratively re-evaluate and refine its generated outputs, ensuring that the final visual state in the multimodal reasoning trajectory achieves better alignment with the prompt and improved visual coherence.
Table 3: Quantitative comparisons on GEdit-Bench [38]. All metrics are reported as higher-is-better ($\uparrow$).
[…]
G_SC G_PQ G_O G_SC G_PQ G_O
Private Gemini 2.0 [33] 6.73 6.61 6.32 5.43 6.78 5.36
GPT-4o [31] 7.85 7.62 7.53 7.67 7.56 7.30
Open-source Instruct-Pix2Pix [6] 3.58 5.49 3.68 - - -
MagicBrush [83] 4.68 5.66 4.52 - - -
AnyEdit [81] 3.18 5.82 3.21 - - -
OmniGen [73] 5.96 5.89 5.06 - - -
Step1X-Edit [39] 7.09 6.76 6.70 7.20 6.87 6.86
BAGEL [15] 7.36 6.83 6.52 7.34 6.85 6.50
Uni-CoT 7.91 6.24 6.74 8.01 6.30 6.87
Table 4: Quantitative comparisons on KRIS [71]. Uni-CoT achieves the top-1 performance among open-source models on the KRIS benchmark and even surpasses the commercial model Gemini 2.0 by 5.59 points in overall score.
Model Perception Conceptual Reasoning Procedural Knowledge Overall
Attribute Spatial Temporal Average Social Natural Average Logical Instruction Average Score
Perception Perception Perception Score Science Science Score Reasoning Decompose Score
Gemini-2.0 [33] 66.33 63.33 63.92 65.26 68.19 56.94 59.65 54.13 71.67 62.90 62.41
[…]
Doubao [7] 70.92 59.17 40.58 63.30 65.50 61.19 62.23 47.75 60.58 54.17 60.70
BAGEL [15] 64.27 62.42 42.45 60.26 55.40 56.01 55.86 52.54 50.56 51.69 56.21
BAGEL-Think [15] 67.42 68.33 58.67 66.18 63.55 61.40 61.92 48.12 50.22 49.02 60.18
Uni-CoT 72.76 72.87 67.10 71.85 70.81 66.00 67.16 53.43 73.93 63.68 68.00
GPT-4o (OpenAI) [31] 83.17 79.08 68.25 79.80 85.50 80.06 81.37 71.56 85.08 78.32 80.09
Table 5: Quantitative comparisons on RISE [85]. Uni-CoT superpass Bagel and show comparable performance with Gemini-2.0.
Model Temporal Causal Spatial Logical Overall
Gemini-2.0 [33] 8.2 15.5 23.0 4.7 13.3
BAGEL-Think [15] 5.9 17.8 21.0 1.2 11.9
BAGEL [15] 2.4 5.6 14.0 1.2 6.1
Uni-CoT 8.2 18.9 20.0 1.2 12.5
GPT-4o [31] 34.1 32.2 37.0 10.6 28.9
Figure 5: Qualitative Results for Reliable Image Editing. Uni-CoT demonstrates considerable image editing abilities, further supporting the effectiveness of its micro-level CoT reasoning. It can generate textual editing instructions and modify the current visual state accordingly.
4.4 Results for Image Editing
Quantitative Results.
First, we evaluate the basic editing capability of Uni-CoT on GEdit-Bench [38]. As shown in Table 3, our model achieves competitive performance compared with other approaches.
Furthermore, we report KRIS and RISE benchmark results in Table 4 and Table 5. Surprisingly, on KRIS benchmark, Uni-CoT outperforms all open-source baselines across perception, conceptual, and procedural categories. Remarkably, it also surpasses the closed-source Gemini 2.0 in overall score, highlighting its robust and interpretable editing capabilities under complex reasoning instructions. On RISE benchmark, Uni-CoT demonstrate comparable performance with Gemini 2.0 with respect to both overall performance on the four reasoning categories and the sub-dimension evaluation metrics of instruction reasoning, appearance consistency and visual plausibility.
Figure 6: Qualitative Results for Nano-Banana like Generation.
Qualitative Analysis.
Figure 5 showcases editing results on challenging instructions. Uni-CoT demonstrates strong fidelity to the original context while making precise visual modifications. As exhibited in this figure, our method achieves high prompt consistency, significant spatial accuracy, and plausible visual transitions, highlighting the effectiveness and interpretability of our CoT-guided editing strategy.
4.5 Nano-banana like Generation
Recently, Google introduced Nano-Banana [44], a unified image generator that achieves results once considered unattainable, such as synthesizing realistic landscapes from satellite imagery or isohypse line maps (top of Figure 6).
We hypothesize that this ability stems from an inherent reasoning-driven generation mechanism. Concretely, the transformation from isohypse maps to landscape images can be decomposed into three steps: (1) 2d-to-3d: interpolating the 2D isohypse map into a 3D terrain; (2) 3d-crop: focusing on a camera view from the 3D representation; and (3) 3d-to-real: rendering the cropped view into a realistic landscape.
To validate this hypothesis, we construct a dataset of 3K multi-modal Chain-of-Thought (CoT) geography samples following this decomposition and refine our model on it. The model converges efficiently and exhibits strong performance on this test set (bottom of Figure 6). In contrast, direct fine-tuning on raw isohypse–landscape pairs results in unstable training and incoherent generations. More results are shown in Figure S3.
5 Conclusion
We present Uni-CoT, a unified Chain-of-Thought framework that enables coherent and grounded multimodal reasoning across vision and language within a single model. By introducing a two-level hierarchical reasoning architecture, comprising a Macro-Level CoT for high-level planning and a Micro-Level CoT for subtask execution modeled as an MDP, we significantly reduce computational complexity and improve reasoning efficiency. Our structured training paradigm further enables effective supervision and preference-based fine-tuning. Extensive experiments across reasoning-driven image generation and editing benchmarks demonstrate Uni-CoT’s superiority in both performance and interpretability. We believe Uni-CoT offers a scalable foundation for future multi-modal reasoning systems.
6 Limitations and Path Forward
In the current implementation of the Uni-CoT framework, we have successfully integrated two core mechanisms that significantly enhance multimodal reasoning capabilities: the Sequential Decomposition Mechanism at the Macro-Level CoT and the Self-Reflection Mechanism at the Micro-Level CoT. These components have yielded notable improvements in reasoning-driven image generation and editing tasks. We will release the corresponding codebase shortly and warmly welcome the community to explore and build upon it.
Despite this progress, several challenges remain. First, we are actively developing more advanced strategies at the Macro-Level CoT, namely the Parallel Decomposition Mechanism and Implicit Planning via Progressive Refinement. While promising, these mechanisms introduce significantly more complex path and memory management, making it difficult to achieve stable training and robust generalization. This is particularly problematic in out-of-distribution scenarios, where reasoning trajectories often deviate from training-time patterns.
Second, our current focus has been on generation tasks, which are more tolerant of imprecise visual state transitions. In contrast, understanding tasks, such as adding auxiliary lines in geometric problems, require strict, step-by-step visual coherence and precise structural alignment. We have observed that directly applying preference learning in such cases leads to suboptimal results, largely due to its inability to enforce fine-grained visual consistency.
To address these limitations, we are exploring new strategies aimed at improving trajectory modeling, memory control, and visual transition accuracy. These include architectural improvements and learning paradigms specifically designed to support physically grounded and logically consistent visual reasoning. These enhancements are currently under development and will be featured in a future release.
We are encouraged by the progress thus far and look forward to sharing more comprehensive results and expanded capabilities in the next version of Uni-CoT.
7 Acknowledgment
We would like to thank Zhengbo Zhang, Qi Lv, Jun Gao and Zhiheng Li for helpful discussion and feedback on this work. The computations in this research were performed using the CFFF platform of Fudan University and QiZhi Platform [1].
Supplementary Material
Appendix A Details of Attention Mask
Figure S1: Attention mask: (a) Macor attention mask; (b) Micro attention mask.
As shown in Figure S1(a), under the Macro masked attention scheme, the model has access only to the system prompt, high-level planning outputs, and final subtask outcomes, while intermediate reasoning traces, such as textual rationales or image edits generated during subtask execution, are completely masked. In contrast, Figure S1(b) illustrates the Micro masked attention scheme, where each self-reflection step can attend only to the immediately preceding state (e.g., the last image-text pair) and the current subtask instruction.
Appendix B Details of Training Paradigm
B.1 Supervised Fine-tuning
We decompose Uni-CoT training into two reasoning levels: Macro-Level CoT learning, which captures global planning and final result synthesis, and Micro-Level CoT learning, which focuses on subtask execution and self-reflection.
Macro-Level CoT Learning.
The macro-level reasoning process consists of two components: global planning and final result synthesis. In the global planning stage, the task is framed as an understanding problem, where the model consumes multi-modal inputs and produces a structured textual plan, trained under a cross-entropy loss $\mathcal{L}_{\text{CE}}^{\text{txt}}$. In the final result synthesis stage, the task is cast as a generative problem: the model integrates both the intermediate multi-modal outputs and the global plan to generate the final solution. For understanding tasks, this corresponds to producing a textual conclusion (supervised by cross-entropy loss $\mathcal{L}_{\text{CE}}^{\text{txt}}$), whereas for image generation tasks, it corresponds to synthesizing both textual and visual outputs (supervised joint loss $\mathcal{L}_{\text{joint}}$ that combines cross-entropy and mean squared error losses). The data structures underlying these two stages are summarized in Table S3.
For convenience, we review the definitions of each loss function mentioned in the main paper below. As discussed in Sec. 2, the cross-entropy loss for textual supervision is defined as:
$\mathcal{L}_{\text{CE}}^{\text{txt}}=-\sum_{i=1}^{C}x_{i}\log(\hat{x}_{i}),$ (1)
where $x_{i}$ denotes the ground-truth token, $\hat{x}_{i}$ represents the predicted probability of that token, and $C$ is the vocabulary size.
Similarly, following Sec. 2, we define the mean squared error (MSE) loss for rectified flow supervision as:
$\mathcal{L}_{\text{MSE}}^{\text{img}}=\mathbb{E}\left[\left\|\mathbf{g}_{\theta}(\mathbf{x}_{t}\mid\mathbf{c})-\left(\mathbf{x}_{0}-\mathbf{x}_{1}\right)\right\|^{2}\right],$ (2)
where $\mathbf{g}_{\theta}$ denotes our model, $\mathbf{x}_{t}$ is a noisy latent obtained by interpolating between a clean latent $\mathbf{x}_{0}$ and a Gaussian noise sample $\mathbf{x}_{1}$, $\mathbf{c}$ is the set of text and image conditions, and $\mathbf{g}_{\theta}(\mathbf{x}_{t}\mid\mathbf{c})$ denotes the predicted velocity vector.
Then the joint loss $\mathcal{L}_{\text{joint}}$ is defined as:
$\mathcal{L}_{\text{joint}}=\lambda_{\text{CE}}\cdot\mathcal{L}_{\text{CE}}^{\text{txt}}+\mathcal{L}_{\text{MSE}}^{\text{img}},$ (3)
where ${\lambda}_{\text{CE}}$ is a coefficient to balance two losses.
Table S3: Details of Macro-Level CoT tasks.
Objective Data Structure Loss
Global Planning [System Prompt, Multi-Modal Inputs, "Planning Prompt"] Cross-Entropy Loss ($\mathcal{L}_{\text{CE}}^{\text{txt}}$)
Result Synthesis [System Prompt, Multi-Modal Inputs, Generated Plan, Multi-Modal Intermediate data, "Final Results"] Joint Loss ($\mathcal{L}_{\text{joint}}$)
The Micro-Level CoT learning.
The micro-level reasoning process consists of two complementary components: subtask completion and self-reflection modeling.
For subtask completion, the model is supervised with interleaved multi-modal signals. Specifically, analogous to the final result synthesis, we employ a joint loss $\mathcal{L}_{\text{joint}}$ that integrates cross-entropy (CE) for textual outputs and mean-squared error (MSE) for text and image predictions, thereby ensuring alignment across modalities.
For self-reflection modeling, as detailed in Sec. 3.2, we formulate the self-reflective procedure within a Markov Decision Process (MDP) framework. Based on this formulation, we introduce four auxiliary objectives for self-reflection modeling: (1) Text Action $a_{t}^{\text{txt}}$ Generation, where the model evaluates the intermediate results and predicts an editing instruction $a_{t}^{\text{txt}}$, supervised by cross-entropy loss $\mathcal{L}_{\text{CE}}^{\text{txt}}$; (2) Image Action $a_{t}^{\text{img}}$ Generation, where the model generates visual modifications $a_{t}^{\text{img}}$ conditioned on the textual instruction $a_{t}^{\text{txt}}$, supervised by mean squared error loss $\mathcal{L}_{\text{MSE}}^{\text{img}}$; (3) Next-State $s_{t+1}$ Prediction, where the model analyzes and summarizes the status of the modified image, supervised by cross-entropy loss $\mathcal{L}_{\text{CE}}^{\text{txt}}$; (4) Reward $r_{t}$ Estimation, where the model regresses to a scalar feedback signal or textual description that measures the quality of the intermediate reasoning step relative to the final task objective, supervised by cross-entropy loss $\mathcal{L}_{\text{CE}}^{\text{txt}}$. The data structure is detailed in Table S3.
Table S3: Details of Micro-Level CoT tasks.
Objective Data Structure Loss
Subtask Completion [System Prompt, Subtask Prompt, "Initial Image & Text Results"] Joint Loss ($\mathcal{L}_{\text{joint}}$)
Text Action $a_{t}^{\text{txt}}$ [System Prompt, Subtask Prompt, Current Image & Text, "Editing Prompt"] Cross-Entropy Loss ($\mathcal{L}_{\text{CE}}^{\text{txt}}$)
Image Action $a_{t}^{\text{img}}$ [System Prompt, Subtask Prompt, Current Image & Text, Editing Prompt, "Edited Image"] Mean Square Error Loss ($\mathcal{L}^{img}_{\text{MSE}}$)
Next-State $s_{t+1}$ [System Prompt, Subtask Prompt, Edited Image, "Image Analysis"] Cross-Entropy Loss ($\mathcal{L}_{\text{CE}}^{\text{txt}}$)
Reward $r_{t}$ [System Prompt, Subtask Prompt, Edited Image, Image Analysis, "Evaluation"] Cross-Entropy Loss ($\mathcal{L}_{\text{CE}}^{\text{txt}}$)
B.2 Reinforcement Learning
We will release the details of Reinforcement Learning Framework in the future.
Figure S2: Data curation pipeline for hierarchical reasoning. We collect T2I prompts from various datasets, then for (a)Macro-level reasoning data pipeline: prompts are first enhanced via GPT-4o or Qwen3, then decomposed into several sequential or parallel subtasks. Then GPT-4o as well as Bagel-Think are used for subtask image generation, evaluation and refinement.(b)Micro-level reasoning data pipeline: we use Bagel-Think model to generate preliminary images based on collected T2I prompts. We then perform several rounds of self-reflection, using GPT-4o to first evaluate on current images and generate refinement instruction, then generate edited images conditioned on refinement instruction.
Appendix C Details of Dataset.
Data curation process.
Figure S2 illustrates the data pipeline we used for data curation. We collect interleaved text-image data for Macro-level and Micro-level reasoning paradigm respectively. We prepare text-to-image generation prompts from multiple datasets as seed prompts for prompt expansion.
For Macro-level reasoning data, we then enhance the prompts in the following two aspects: (1) We first check if the given prompts entail domain expertise knowledge or common-sense reasoning. If such reasoning or logical induction exist in prompts, we rewrite the prompts to explicitly state the result of reasoning deduction. For example, if the original prompt describes ”a melting ice cream cone in desert sun”, then the deduced rewritten prompt should elaborate on the object states under given condition, i.e. ”dripping, puddle on hot sand” (2) We then enrich the generation prompts via adding auxiliary visual detail or attributes to concretely illustrate any abstract concepts or description, especially those regarding art styles, vibes, environment, etc. After prompt enhancement, we employ models such as GPT-4o or Qwen-plus to decompose the enhanced prompts into 2-3 subtasks such that complex image generation goal inferred by the enhanced prompts are broken into simpler and logically coherent sequential or parallel sub-goal. We then utilize models capable of image generation such as BAGEL-Think or GPT-4o, to perform image generation and editing following the subtask instruction. We also employ VLM models (GPT-4o) to evaluate on the results of subtask execution and generate corresponding subtask refinement instruction. The textual detail of subtask planning, subtask evaluation and refinement, and the intermediate images generated in all subtask steps, are all collected as interleaved Macro-level data.
For Micro-level reasoning data, we directly generate preliminary images via BAGEL-Think. Next we perform several rounds of self-reflection. We repetitively evaluate on generated or edited images from the previous round using VLM models such as GPT-4o, outputting an assessment of image quality and instruction-following as well as an instruction on how should the image be refined in order to align with the intention of the original prompts better. We then employ image generation models (i.e. GPT-4o) capable of image editing to edit the images in prior rounds according to the refinement instruction. We collect the textual and visual output in each self-reflection loop as the interleaved text-image data for Micro-level reasoning SFT.
Data information.
In total, we initially collect approximately 11K complete long-form multi-modal Chain-of-Thought trajectories, which can be decomposed into 11K samples for macro-level learning and 20K samples for micro-level learning. Moreover, thanks to the decomposable nature of our framework, the model can also benefit from fragmented data. To further enhance its basic capabilities, we additionally incorporate 114K text-to-image generation samples, 68K samples from echo4o dataset and 46K samples from sharegpt4o-image dataset, and 46K image editing samples from sharegpt4o-image. We will detail the data distribution in the future.
Appendix D More Experiments.
Figure S3: More Visualization of our Macro-CoT thinking process.